OpenStudy (anonymous):

A line passes through the point (0, 2) and has a slope of -1/4 What is the equation of the line?

OpenStudy (anonymous):

The equation you'll be using is: \[y-y_1=m(x-x_1)\]Here, m is the slope of the line, and x_1,y_1 are the x and y coordinates of any point on the line. You have the slope and a point, so you can just plug in and simplify as needed.

OpenStudy (anonymous):

so, how should it look like?

OpenStudy (anonymous):

?

OpenStudy (jdoe0001):

point-slope form -> \(\huge y-y_1=m(x-x_1)\) m = slope (0, 2) = \(\bf (x_1, y_1)\)

OpenStudy (jdoe0001):

plug your values in, solve for "y" to get your equation

OpenStudy (anonymous):

so how will that look

OpenStudy (jdoe0001):

well, plug in the values, and you'd know

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

yeah im lost

OpenStudy (jdoe0001):

well, did you plug in the values? the point-slope form takes, slope, and "x" and "y" coordinates for a point given you're given all 3 of those

OpenStudy (anonymous):

4x+y+2?

OpenStudy (jdoe0001):

hmmm, the slope given is -1/4